OpenStudy (anonymous):

given f(x)=2(x+1)^2+3. Find the greatest value of 1/f(x)

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

how to do that?? why not 3??

OpenStudy (anonymous):

as the smallest value of f(x) is 3, the greatest value of 1/f(x) is 1/3

OpenStudy (anonymous):

the least value inside the square is zero so the least value of f(x)= 2* 0+3= 3,

OpenStudy (anonymous):

my idea is: g(x) = 1/f(x) then g'(x) = -f'(x)/(f(x))^2

OpenStudy (anonymous):

saruz's method is much easier =D

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

see u r trying to find out the maximum value of \[\frac{1}{f(x)}\]so the minimum value of f(x) will give u the maximum vaue for the \[\frac{1}{f(x)} = \frac{1}{2(-1 + 1)^2 + 3} = \frac{1}{(2 \times 0) + 3} = \frac{1}{3}\]